### v2

# Defending Jailbreak Attack in VLMs via Cross-modality Information Detector

Abstract
Vision Language Models (VLMs) extend the capacity of LLMs to comprehensively understand vision information, achieving remarkable performance in many vision-centric tasks. Despite that, recent studies have shown that these models are susceptible to jailbreak attacks, which refer to an exploitative technique where malicious users can break the safety alignment of the target model and generate misleading and harmful answers. This potential threat is caused by both the inherent vulnerabilities of LLM and the larger attack scope introduced by vision input. To enhance the security of VLMs against jailbreak attacks, researchers have developed various defense techniques. However, these methods either require modifications to the model’s internal structure or demand significant computational resources during the inference phase. Multimodal information is a double-edged sword. While it increases the risk of attacks, it also provides additional data that can enhance safeguards. Inspired by this, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, utilizing the cross-modal similarity between harmful queries and adversarial images. This simple yet effective cross-modality information detector, CIDER, is independent of the target VLMs and requires less computation cost. Extensive experimental results demonstrate the effectiveness and efficiency of CIDER, as well as its transferability to both white-box and black-box VLMs.
Defending Jailbreak Attack in VLMs via Cross-modality Information Detector
Yue Xu1thanks: equal contribution , Xiuyuan Qi1* , Zhan Qin2 , Wenjie Wang1 thanks: W.Wang is the corresponding author. 1School of Information Science and Technology, ShanghaiTech University 2College of Computer Science and Technology, Zhejiang University {xuyue2022,qixy1,wangwj1}@shanghaitech.edu.cn, qinzhan@zju.edu.cn
1 Introduction
The remarkable advancements in Large Language Models (LLMs) have significantly improved performance benchmarks in various natural language processing (NLP) tasks Achiam et al. (2023); Touvron et al. (2023); Zhao et al. (2023); Chiang et al. (2023). To extend the capacities and open up the potentials of LLMs in comprehensively understanding diverse types of data, such as visual information, researchers have developed Vision Language Models (VLMs) that integrate visual modalities to handle vision-centric tasks. VLMs use LLMs as a core, complemented by modal-specific encoders and projectors, enabling them to process, reason, and generate outputs from multimodal data Yin et al. (2023); Dai et al. (2024); Bai et al. (2023). A typical VLM architecture is illustrated in Figure 1.
Figure 1: The illustration of a typical VLM architecture.
The widespread adoption of VLMs in various applications brings significant safety challenges, particularly due to inherited vulnerabilities from traditional LLMs, such as the susceptibility to jailbreak attacks Carlini et al. (2024); Li et al. (2024); Qi et al. (2024). Jailbreak attacks refer to an exploitative technique where malicious users can craft sophisticated-designed prompts to lead LLMs to answer misleading or harmful questions, effectively breaking the alignment and bypassing the model’s safeguard. Various jailbreak attack algorithms targeting LLMs have been proposed, which can be categorized into template-based Deng et al. (2024); Chao et al. (2023); Li et al. (2023) and optimize-based Zou et al. (2023) approaches.
Additionally, VLMs not only inherit the vulnerabilities of LLMs but also become more susceptible to jailbreak attacks due to their integration of the visual modality. On the one hand, jailbreak attacks against VLMs can originate from both the textual and visual modalities, significantly broadening the scope of potential adversarial examples Gong et al. (2023); Shayegani et al. (2023). On the other hand, recent research indicates that fine-tuning VLMs to learn the vision modality can cause LLMs to disregard their previously learned safety alignment Zong et al. (2024).
Figure 2: The workflow of safeguarding VLM against jailbreak attacks via CIDER.
The existing jailbreak attacks on VLMs can be categorized into two strategies. One is white-box optimization-based attacks, which define a loss function to generate imperceptible perturbations in the image modality Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024). The other is black-box strategies including typographically transforming harmful queries into images such as FigStep Gong et al. (2023) or adding related images containing harmful text such as QR Liu et al. (2023).
From the defense perspective, optical character recognition (OCR) can serve as an effective detection tool for the second strategy but fails when defending against optimization-based adversarial attacks. In addition, Zong et al. (2024) creates a vision-language dataset named VLGuard containing both safe and unsafe queries and images, which can be used to fine-tune VLMs for improved safety against jailbreak attacks. However, the effectiveness of VLGuard is only tested on FigStep attack and it requires the model to be white-box to fine-tune. Zhang et al. (2023) proposed a mutation-based jailbreaking detection framework named Jailguard. However, the performance of Jailguard heavily relies on the VLMs’ original safety alignment, and it significantly increases computational costs during the inference phase.
Multimodal information is a double-edged sword: while it increases the risk of attacks, it also provides additional data that helps enhance safeguards. Inspired by this potential, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, specifically targeting optimization-based jailbreak attacks that are more imperceptible and susceptible. The intuition is that optimization-based perturbations break the VLM’s safeguards by capturing the main harmful content in the malicious query. As a result, the semantic distance between a harmful query and an adversarially perturbed image is significantly smaller than that between a harmful query and a clean image.
Directly utilizing the difference between clean and adversarial images on the semantic distance to harmful query is challenging, as the absolute value of the distance varies across different harmful queries. To address this issue, we incorporate a denoiser to preprocess the vision modality, using the relative shift in the semantic distance before and after denoising to reflect the difference between clean and adversarial images. As shown in Figure 2, the key insight of CIDER is to identify whether an image is adversarially perturbed based on the semantic similarity between image and text modality before and after denoising ($\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle-\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle$). If the image modality is not perturbed, the semantic similarity between text and image remains stable. However, the adversarially perturbed image designed for jailbreak will experience a significant drop. By setting a threshold based on this change, we can effectively detect adversarially perturbed images aimed at jailbreaking VLMs. The detailed intuition is elaborated in Section 2.
As a pre-detection module encapsulated before any VLMs, the key advantage of CIDER is its plug-and-play nature, making it independent of the target model. Additionally, timely inference is crucial for safeguarding VLMs. CIDER achieves this by adding only denoiser procedures, ensuring efficient without introducing significant inference latency.
In this work, we propose CIDER, an effective and efficient pre-detection module that denoises and inspects each input image. For images identified as adversarially perturbed for jailbreak purposes (where the semantic shift exceeds a pre-defined threshold), the VLM will refuse to generate a response. Images deemed normal will be processed along with the text input for model inference by the VLM. The workflow of safeguarding VLMs against jailbreak attacks using CIDER is illustrated in Figure 2. Our contribution can be summarized as follows:
  • Based on the intuition that cross-modality information is a double-edged sword, we investigate the relationship between malicious queries and adversarial perturbed images in the semantic space. By incorporating a diffusion-based denoiser to uncover the potential of mitigating harmful information in adversarial images through denoising.
  • We propose a plug-and-play jailbreaking detector, CIDER, which can effectively safeguard VLMs while incurring almost no additional computational overhead.
  • Extensive experiments validate that CIDER outperforms the baseline method, achieving a higher detection success rate while reducing the computational cost as well. Experimental results also demonstrate its strong transferability across both white-box and black-box VLMs and attack methods.
2 Intuition: Cross-modality information is a double-edged sword
While multimodal information aggravates model vulnerability to jailbreak attacks, it also provides additional information for defense. The design of CIDER is based on the intuition that optimization-based jailbreak attacks break the VLM’s safeguards by sharing harmful content in the malicious query to the image modality. Consequently, the adversarially perturbed image is closer to the harmful query in the semantic space than the clean images. To support this intuition, we first explain the fundamentals of the optimization-based jailbreak attacks on VLMs. Then, we design a few experiments to explore how cross-modal analysis can help safeguard VLMs, and we analyze the semantic difference between clean and adversarial images relative to harmful queries, both before and after denoising.
2.1 Preliminaries: Optimization-based Jailbreak Attacks on VLMs
Optimization-based VLM jailbreaking is similar to adversarial attacks on image classification tasks Goodfellow et al. (2014), with the primary difference being the difference in the loss function. Specifically, given a dataset $D=\{(q,a)\}$ where $q$ represents the harmful queries and $a$ is the corresponding targeted answers, the attacker aims to find an adversarial image $x_{adv}$ that can encourage the VLM $\mathcal{F}$ to generate $a$ when inputting $q$ along with $x_{adv}$. The objective can be formulated as:
$x_{adv}=\underset{x_{adv}\in[0,1]^{d}}{\mathrm{argmin}}log(\mathcal{F}(a|q,x_{adv}))\vspace{-.5em}$ (1)
where $\mathcal{F}(a|q,x_{adv})$ represents the likelihood that the VLM $\mathcal{F}$ generate answer $a$ when given the adversarial image $x_{adv}$ and the query $q$.
2.2 Experimental Setup
We design a series of experiments to explore how cross-modality information can help safeguard VLMs and to analyze the semantic difference between clean and adversarial images to harmful queries, before and after denoising. We utilize the image and text encoder of the state-of-the-art VLM LLaVA-v1.5-7B Liu et al. (2024) to capture the semantic meanings. To measure the semantic similarity, we employed cosine similarity which is a standard metric widely used in information retrieval and natural language processing Park et al. (2020); Pal et al. (2021). In terms of denoiser, we incorporate a diffusion-based denoiser Nichol and Dhariwal (2021) to preprocess the image modality.
The inputs to the VLMs consist of two modalities: images and text queries. For malicious queries, we utilize the validation set proposed in the Harmbench framework Mazeika et al. (2024), which contains 40 textual harmful behaviors across 7 semantic categories. For images, we use 5 adversarial images generated by an optimization-based jailbreak attack Qi et al. (2024) and 5 clean images from ImageNet Deng et al. (2009). As a result, we have 200 adversarial text-image pairs and 200 clean pairs.
2.3 Findings
According to the results displayed in Figure 3, the key findings can be summarized as follows:
(a) $\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E^{\textit{C}}_{\textit{img(o)}}}\rangle-\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E^{\textit{A}}_{\textit{img(o)}}}\rangle$
(b) $\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle$
(c) $\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle$
(d) $\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle-\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle$
Figure 3: Experimental result. (a) the distribution of the difference between clean and adversarial images regarding their cos-sim with harmful queries. (b) the distribution of cos-sim between harmful queries and clean/adversarial images. (c) the change of the cos-sim during denoising. (d) the distribution of $\Delta$cos-sim before and after denoising of clean/adversarial images.
Finding 1: Adversarial images indeed contain harmful information.
For each harmful query, we calculate the cosine similarity between the queries and both clean and adversarial images, denoted as $\langle\bm{E^{M}_{\textit{text}}},\bm{E^{C}_{\textit{img(o)}}}\rangle$ and $\langle\bm{E^{M}_{\textit{text}}},\bm{E^{A}_{\textit{img(o)}}}\rangle$ respectively. Figure 3(a) shows the distribution of $\langle\bm{E^{M}_{\textit{text}}},\bm{E^{C}_{\textit{img(o)}}}\rangle-\langle\bm{E^{M}_{\textit{text}}},\bm{E^{A}_{\textit{img(o)}}}\rangle$. It can be observed that the distribution is almost entirely concentrated in the negative region, indicating that, for a specific harmful query, the semantic distance between it and an adversarial image is smaller than that between it and a clean image. Therefore, we can conclude that adversarial images indeed carry harmful information from queries.
Finding 2: Directly utilizing the semantic difference between clean and adversarial images to harmful query is challenging
Figure 3(b) shows the distribution of the absolute value of $\langle\bm{E^{M}_{\textit{text}}},\bm{E^{C}_{\textit{img(o)}}}\rangle$ and $\langle\bm{E^{M}_{\textit{text}}},\bm{E^{A}_{\textit{img(o)}}}\rangle$. Although the distribution differs in the peak and concentration, distinguishing between adversarial and clean images based solely on the absolute value of the difference is challenging. This difficulty arises because the cosine similarity between different queries and adversarial images varies significantly, and the absolute value of the difference does not fully capture the characteristics of the images.
Finding 3: Denoising can reduce harmful information but cannot eliminate
Subsequently, we applied denoising to each image 350 times, assessing cosine similarity with harmful queries every 50 iterations (visualization of the denoising is relegated to Appendix A). Figure 3(c) illustrates how cosine similarity between harmful query and adversarial images decreases as denoising progresses, indicating a reduction in harmful information. Despite this reduction, denoised adversarial images, when tested with harmful text inputs in the VLM, still enabled a significant number of queries to jailbreak. Thus, while denoising mitigates harmful information in images, it does not eliminate their adversarial properties.
Finding 4: Relative shift in the semantic distance before and after denoising can help detect adversarial images.
In Figure 3(d), we present the change of cosine similarity before and after denoising ($\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle-\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle$). The distribution of cosine similarity between harmful queries and images shifts notably after denoising, contrasting with the distribution in Figure 3(b). This observed shift supports our intuition that relative changes before and after denoising, rather than absolute differences, provide a method to distinguish adversarial images. Figure 3(c) further demonstrates this trend, showing a consistent decrease in cosine similarity between adversarial images and harmful queries, while the cosine similarity with clean images varies less.
3 Method
In this section, we first overview the defense pipeline and the components of CIDER, followed by introducing the threshold selection strategy.
3.1 Overview
CIDER is implemented on top of the VLMs to defense optimization-based adversarial jailbreak attacks. Figure 2 presents the overview of the CIDER pipeline. Specifically, given a text-image input pair, denoted as <text, img(o)>, CIDER calculates the embeddings of text and image modalities, denoted as $\bm{E_{\textit{text}}}$ and $\bm{E_{\textit{img(o)}}}$. Then, the built-in denoiser in CIDER will perform 350 denoising iterations on the image(o), calculating the denoised embeddings every 50 iterations, denoted as $\mathcal{E}=\bm{E_{\textit{img(d)}}}$. The img(o) will be identified as an adversarial example if any $\bm{E_{\textit{img(d)}}}\in\mathcal{E}$ satisfy the following condition:
$\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle-\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle>\tau$ (2)
[…]
Input: $img(o)$: input image, $text$: input query, $\mathcal{F}$: target VLM, $\tau$: predefined threshold.
$flag\leftarrow true$;
for $i\leftarrow 0$ to $350$ Step $50$ do
       $img(d)\leftarrow denoiser(img(o),i)$;
       $\bm{E_{\textit{text}}}\leftarrow TextEncoder(text)$;
       $\bm{E_{\textit{img(o)}}}\leftarrow ImgEncoder(img(o))$;
       $\bm{E_{\textit{img(d)}}}\leftarrow ImgEncoder(img(d))$;
       $d\leftarrow\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle-\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle$;
       if $d>\tau$ then
             $flag\leftarrow false$;
if $flag=true$ then
       Return $\mathcal{F}(img(o),text)$;
else
       Return "I’m sorry, but I can not assist."
Algorithm 1 CIDER defense pipeline
3.2 Threshold selection
The threshold is selected based on the harmful queries and clean images ensuring that the vast majority of clean images pass the detection. The selection of threshold $\tau$ can be formulated as:
$r=\frac{\sum\mathbb{I}(\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E^{\textit{C}}_{\textit{img(o)}}}\rangle-\langle\bm{E^{\textit{M}}_{\textit{text}}},\bm{E^{\textit{C}}_{\textit{img(d)}}}\rangle<\tau)}{\#\textit{samples}}$ (3)
where $r$ represents the passing rate and $\bm{E^{\textit{M}}_{\textit{text}}}$, $\bm{E^{\textit{C}}_{\textit{img(o)}}}$, $\bm{E^{\textit{C}}_{\textit{img(d)}}}$ stand for the embeddings of input query, the input image and denoised image respectively. The threshold $\tau$ is determined by controlling the passing rate $r$. For example, using the $\tau$ when setting $r$ to 95% as the threshold indicates allowing 95% percent of clean images to pass the detection.
The selection of the threshold significantly impacts the effectiveness of CIDER: a threshold that is too high will cause many adversarial examples to be classified as clean samples, resulting in a low true positive rate (TPR); conversely, a threshold that is too low will lead to a high false positive rate (FPR), affecting the model’s normal performance.
Figure 4: TPR-FPR trade-off on validation set.
The ablation study is conducted to determine the optimal threshold. By treating adversarial pairs as positive samples and clean pairs as negative samples, we plot the TPR-FPR curve with thresholds ranging from 80% to 100% in 1% increments, as shown in Figure 4. Ideally, we expect high TPR and low FPR (the upper left corner of the plot). Therefore, we selected $\tau$ when $r$ equals 95% as the detection threshold of CIDER.
4 Experiment
In this section, we begin by outlining the configurations of our experiments, including the models, datasets, baselines, and evaluation metrics. We then evaluate the effectiveness and efficiency of CIDER, comparing with the baseline methods. Next, we discuss the trade-off between robustness and utility. Finally, we demonstrate the generalization of our method.
4.1 Configurations
Models. Note that CIDER is an auxiliary model that is independent to the VLMs. We use LLaVA to capture the semantic meaning of each modality, but CIDER can be plugged into any other VLMs. To demonstrate the effectiveness of CIDER, we test the detection and defense performance on four open-source VLMs, including LLaVA-v1.5-7B Liu et al. (2024), MiniGPT4 Zhu et al. (2023), InstructBLIP Dai et al. (2024), and Qwen-VL Bai et al. (2023), as well as the API-access VLM, GPT4V Achiam et al. (2023).
Figure 5: ASR of base VLM, defending with CIDER and defending with Jailguard
Datasets. Similar to the dataset used in Section 2.2, we generate 800 adversarial text-image pairs utilizing the 160 harmful queries in Harmbench Mazeika et al. (2024) and adversarial images provided by Qi et al. (2024). To further demonstrate CIDER will not destroy the original utilities on the normal queries, we also evaluate the utility of CIDER protected VLMs on MM-Vet benchmark Yu et al. (2023), which examines 6 core vision language capabilities, including recognition, optical character recognition (OCR), knowledge, language generation, spatial awareness, and math.
Baseline and evaluation metrics. We use Jailguard Zhang et al. (2023) as a baseline, which is a SoTA mutation-based jailbreak detection strategy that protects the VLMs at the inference stage. We involve four evaluation metrics to demonstrate the performance of defending methods from different aspects. From the perspective of the effectiveness of CIDER, we incorporate detection success rate (DSR) and Attack success rate (ASR). DSR represents the proportion of adversarial examples $\mathcal{D}$ that can be successfully detected:
[…]
ASR is a standard evaluation metric indicating the proportion of samples that can successfully jailbreak VLM $\mathcal{F}$ and generate harmful contents, which can be stated as:
[…]
$\mathcal{G}$ refers to an LLM classifier Mazeika et al. (2024) that determines the harmfulness of a response. $\mathbb{I}_{harm}$ and $\mathbb{I}_{harm}$ represent the adversarial and harmful indicator. In terms of efficiency, we measure the time cost of processing a single query. In addition, to evaluate the model utility on regular tasks, responses, we incorporate an online evaluator MM-Vet-Evaluator (2024) provided along with MM-Vet benchmark, which utilizes GPT-4 to generate a soft grading score from 0 to 1 for each answer.
4.2 Effectiveness
DSR. We first demonstrate the overall DSR that CIDER can achieve and compare it with the baseline method, Jailguard. Table 1 shows that CIDER achieves a DSR of approximately 80%, while the DSR of Jailguard varies, depending on the target VLMs. Note that CIDER is independent of the VLMs, thus the DSR does not vary with the choice of VLMs. However, Jailguard’s detection capability relies heavily on the model’s safety alignment, so the DSR also varies. VLMs with good alignment achieve high DSR (e.g., GPT4V), while poorly aligned VLMs have relatively low DSR (e.g., InstructBLIP). In other words, Jailguard does not significantly enhance VLM robustness against adversarial jailbreak attacks, whereas CIDER does. Nonetheless, CIDER achieves a higher DSR than most of the Jailguard results, except Jailguard on GPT4V.
Method detection success rate ($\uparrow$)
Jailguard with LLaVA-v1.5-7B $39.50\%$
Jailguard with InstructBLIP $32.25\%$
Jailguard with MiniGPT4 $69.50\%$
Jailguard with Qwen-VL $77.50\%$
Jailguard with GPT4V $94.00\%$
CIDER $79.69\%$
Table 1: DSR of CIDER and Jailguard
ASR. To evaluate the effectiveness of CIDER, we measure the decline in ASR after applying CIDER. Figure 5 compares the original ASR without defense (red bar), ASR after CIDER (blue bar) and ASR after Jailguard (yellow bar). Note that, Jailguard is solely designed to detect jailbreak input. To ensure a fair comparison, we add an output module following Jailguard’s detection. Specifically, if Jailguard detects a jailbreak, it will refuse to respond, similar to CIDER. Otherwise, the original input will be processed by the VLM.
Across all models, defending with CIDER significantly reduces the ASR, yielding better results than the baseline. This indicates that CIDER effectively enhances the robustness of VLMs against optimization-based jailbreak attacks. The most notable improvements are seen in LLaVA-v1.5-7B, where ASR drops from 60% to 0%, and in MiniGPT4, from 57% to 9%. For VLMs with initially low ASRs, such as InstructBLIP and Qwen-VL, ASR is reduced to approximately 2% and 1% respectively. Another notable disadvantage of Jailguard is observed in models like GPT4V, InstructBLIP, and Qwen-VL, which already had strong safety alignment and resistance to adversarial attacks. In these cases, the use of Jailguard resulted in a slight increase in ASR.
4.3 Efficiency
Timely inference is crucial for safeguarding VLMs in real-world applications. Table 2 shows the time required to process a single input pair and generate up to 300 tokens with different VLMs, comparing no defense, CIDER, and Jailguard.
Model Original CIDER Jailguard
LLaVA-v1.5-7B $6.39s$ $7.41s$ (1.13×1.13$\times$1.13 ×) $53.21s$ (8.32×8.32$\times$8.32 ×)
InstructBLIP $5.46s$ $6.48s$ (1.22×1.$2$) $47.83s$ (8.76×8.76$\times$8.76 ×)
MiniGPT4 $37.00s$ $38.02s$ (1.03×1.03$\times$1.03 ×) $313.78s$ (8.48×8.48$\times$8.48 ×)
Qwen-VL $6.02s$ $7.04s$ (1.19×1.19$\times$1.19 ×) $48.48s$ (8.05×8.05$\times$8.05 ×)
GPT4V $7.55s$ $8.57s$ (1.16×1.16$\times$1.16 ×) $61.04s$ (8.08×8.08$\times$8.08 ×)
Table 2: Time cost to process a single pair of inputs.
CIDER surpasses Jailguard in efficiency, adding only 1.02 seconds per input pair on average, which is relatively acceptable compared to the original inference time. In contrast, Jailguard requires 8-9 times the original processing time. Additionally, CIDER detection is irrelevant to the number of generated tokens in the query answers. Therefore, CIDER does not cause additional overhead when increasing the number of generated tokens, ensuring the stability of CIDER’s efficiency.
4.4 Robustness-utility trade-off
Figure 6: VLM performance with and without CIDER on MM-Vet.
To further demonstrate CIDER’s influence on the original utilities on normal queries, we also evaluate the utility of CIDER protected VLMs on MM-Vet benchmark, including recognition, OCR, knowledge, language generation, spatial awareness, and math. As shown in Figure 6, employing CIDER leads to an approximate 30% overall performance decline on normal tasks. Specifically, CIDER mostly affects the VLM’s recognition, knowledge, and language generation capabilities, while it has minimal impact on OCR, spatial awareness, and math skills. We hypothesize that CIDER’s stringent decision-making process, which outright rejects tasks once an image is identified as adversarial, hampers the model’s overall performance. To further illustrate the robustness-utility trade-off, we conducted an ablation study using denoised images as inputs for the adversarial images, termed CIDER-de. The result is relegated to Appendix B.
To find the optimal balance between safety and utility, we could design a more flexible rejection strategy, such as implementing multi-level thresholds for different types of content. This approach could reduce the negative impact on the model’s functionality and we leave it to our future work.
4.5 Generalization
In the previous sections, we evaluated the ASR and DSR against adversarial examples generated by Qi et al. (2024). To assess the generalization of our defense method, which is crucial for its applicability to other attacks, we evaluate CIDER against another optimization-based jailbreak attack. We generated 800 adversarial pairs using ImgJP, as proposed by Niu et al. (2024). Table 3 presents the drop in ASRs on four open-source VLMs. The ASR for all VLMs dropped to below 4%, with Qwen reaching 0%. Additionally, CIDER achieves a DSR of 93.87% against ImgJP. These results demonstrate that CIDER effectively generalizes in defending against optimization-based adversarial attacks, highlighting its practical utility for real-world applications.
Base CIDER
Model ASR(%) ASR(%) $\Delta$ (%)
LLaVA-v1.5-7B $61.0$ $3.5$ $57.5$
InstructBLIP $4.0$ $1.5$ $2.5$
MiniGPT4 $52.5$ $4.0$ $48.5$
Qwen-VL $6.5$ $0.0$ $6.5$
Table 3: Generalization of CIDER to ImgJP
5 Related Work
Vision Language Model. A typical Vision Language Model (VLM) consists of an image encoder Dosovitskiy et al. (2020) to extract feature maps, a projector to align image modality information with text modality, and a Large Language Model (LLM) to integrate textual and visual input for generating responses. The impressive multimodal capabilities of these models have spurred significant research interest, leading to contributions from both academia and industry Achiam et al. (2023); Liu et al. (2024); Zhu et al. (2023); Dai et al. (2024); Bai et al. (2023).
Jailbreaking VLMs. Incorporating visual information into the LLM framework significantly broadens its range of applications but also introduces new security vulnerabilities, complicating the security issues of VLMs. Besides transferring text jailbreak templates from LLMs to VLMs Luo et al. (2024), effective strategies for jailbreaking VLMs include using gradient-based methods to generate adversarial images Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024), and submitting screenshots containing harmful instructions Gong et al. (2023) or related images Liu et al. (2023); Shayegani et al. (2023). This paper focuses on safeguarding VLMs against gradient-based adversarial image attacks, aiming to fortify VLMs against such sophisticated threats and ensure their robustness and reliability in practical applications.
Safeguarding VLMs. Various defense mechanisms have been proposed to address vulnerabilities in VLMs and enhance their security and robustness. These mechanisms can be categorized into proactive and reactive defenses based on their preventive and responsive nature. Proactive defenses aim to prevent attacks through techniques like adversarial training Zong et al. (2024) and reinforcement learning Chen et al. (2023) during the training phase. In contrast, reactive defenses focus on detecting attacks during the inference phase using methods such as Wang et al. (2024a); Pi et al. (2024); Wang et al. (2024b). However, many of these methods require access to internal model parameters or rely on additional large models for implementation. Our approach prioritizes a reactive defense strategy for its practicality and ease of implementation. Notably, Jailguard Zhang et al. (2023) is closely related to our work, as it detects jailbreak queries by analyzing variations in responses to perturbed inputs. However, Jailguard’s detection success heavily depends on the safety of the underlying LLM and involves significant computational costs.
6 Conclusion
In this work, we propose a plug-and-play cross-modality information detector, CIDER, which can effectively and efficiently defend against adversarial jailbreak attacks. Compared to previous methods, CIDER achieves superior defense performance, as evidenced by higher DSR and a significant decline in ASR, while greatly reducing processing time. We also evaluate the transferability of CIDER to other optimization-based adversarial attacks and demonstrate the robustness-utility trade-off in VLMs. In future research, we aim to improve CIDER by reducing the negative impact on VLM utilities to normal tasks. Additionally, it would be useful to develop defense mechanisms against non-optimization-based jailbreak attacks.
Limitations
We outline the limitations of our study as follows:
1. While CIDER is an effective, efficient, and user-friendly defense mechanism, it does impact VLM performance to some extent. We believe this is due to CIDER’s stringent handling of adversarial examples. In future work, we plan to implement multi-level thresholds to process adversarial examples with varying degrees of rigor, aiming to maintain robust defense capabilities without compromising VLM performance.
2. CIDER is specifically designed to defend against optimization-based adversarial jailbreak attacks, and its effectiveness against other types of jailbreak attacks is uncertain. Future research will explore CIDER’s effectiveness against these alternative attacks and develop corresponding defense strategies, aiming to enhance the overall security and resilience of VLMs against a wider array of adversarial threats.
Ethics Statement
Ensuring the security of Vision Large Language Models (VLMs) is crucial as they become more widely used in various applications. This paper introduces CIDER, a simple yet effective cross-modality information detector designed to defend against adversarial jailbreak attacks in VLMs. Our work significantly contributes to the field by providing a tool that mitigates known vulnerabilities and lays the groundwork for future improvements in safety measures. While CIDER marks significant progress, it doesn’t make VLMs immune to all threats. Continuous evaluation and updates are crucial as VLMs evolve. By sharing CIDER and our findings, we aim to encourage ongoing research and collaboration, promoting advanced and secure VLMs. We are committed to addressing the ethical implications of VLM deployment, ensuring transparency, and prioritizing the responsible use of these technologies for societal benefit.
Appendix A Visualization of denoising
Figure 7 presents an example of an adversarially perturbed image, showing the effects of denoising it after 100, 200, and 300 iterations.
Figure 7: An example of the denoising procedure.
Appendix B Ablation study on robustness-utility trade-off
To further illustrate the robustness-utility trade-off, we perform an ablation study using denoised images as inputs for adversarial images, referred to as CIDER-de. Figure 8 shows the ASR of CIDER-de and Figure 9 shows the MM-Vet score of it. It can be observed that using CIDER-de hardly impacts the utility of the VLM. However, this comes at the expense of greatly diminished defensive effectiveness.
Figure 8: ASR of base VLM, defending with CIDER-de, CIDER and Jailguard
Figure 9: MM-Vet score of base VLM, defending with CIDER-de, CIDER and Jailguard

### v4

@@
-Defending Jailbreak Attack in VLMs via Cross-modality Information Detector
+Cross-modality Information Check for Detecting Jailbreaking in Multimodal Large Language Models
 Abstract
-Vision Language Models (VLMs) extend the capacity of LLMs to comprehensively understand vision information, achieving remarkable performance in many vision-centric tasks.
+Multimodal Large Language Models (MLLMs) extend the capacity of LLMs to understand multimodal information comprehensively, achieving remarkable performance in many vision-centric tasks.
 Despite that, recent studies have shown that these models are susceptible to jailbreak attacks, which refer to an exploitative technique where malicious users can break the safety alignment of the target model and generate misleading and harmful answers.
 This potential threat is caused by both the inherent vulnerabilities of LLM and the larger attack scope introduced by vision input.
-To enhance the security of VLMs against jailbreak attacks, researchers have developed various defense techniques.
+To enhance the security of MLLMs against jailbreak attacks, researchers have developed various defense techniques.
 However, these methods either require modifications to the model’s internal structure or demand significant computational resources during the inference phase.
@@
 Inspired by this, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, utilizing the cross-modal similarity between harmful queries and adversarial images.
-This simple yet effective cross-modality information detector, CIDER, is independent of the target VLMs and requires less computation cost.
-Extensive experimental results demonstrate the effectiveness and efficiency of CIDER, as well as its transferability to both white-box and black-box VLMs.
-Defending Jailbreak Attack in VLMs via Cross-modality Information Detector
+CIDER is independent of the target MLLMs and requires less computation cost.
+Extensive experimental results demonstrate the effectiveness and efficiency of CIDER, as well as its transferability to both white-box and black-box MLLMs.
+The resource is available at https://github.com/PandragonXIII/CIDER.
+Cross-modality Information Check for Detecting Jailbreaking in Multimodal Large Language Models
 Yue Xu1thanks: equal contribution , Xiuyuan Qi1* , Zhan Qin2 , Wenjie Wang1 thanks: W.Wang is the corresponding author.
-1School of Information Science and Technology, ShanghaiTech University 2College of Computer Science and Technology, Zhejiang University {xuyue2022,qixy1,wangwj1}@shanghaitech.edu.cn, qinzhan@zju.edu.cn
+1School of Information Science and Technology, ShanghaiTech University 2The State Key Laboratory of Blockchain and Data Security, Zhejiang University {xuyue2022,qixy1,wangwj1}@shanghaitech.edu.cn, qinzhan@zju.edu.cn
 1 Introduction
@@
 (2023).
-To extend the capacities and open up the potentials of LLMs in comprehensively understanding diverse types of data, such as visual information, researchers have developed Vision Language Models (VLMs) that integrate visual modalities to handle vision-centric tasks.
-VLMs use LLMs as a core, complemented by modal-specific encoders and projectors, enabling them to process, reason, and generate outputs from multimodal data Yin et al.
+To extend the capacities and open up the potentials of LLMs in comprehensively understanding diverse types of data, such as visual information, researchers have developed Multimodal Large Language Models (MLLMs) that integrate visual modalities to handle vision-centric tasks.
+MLLMs use LLMs as a core, complemented by modal-specific encoders and projectors, enabling them to process, reason, and generate outputs from multimodal data Yin et al.
 (2023); Dai et al.
@@
 (2023).
-A typical VLM architecture is illustrated in Figure 1.
-Figure 1: The illustration of a typical VLM architecture.
-The widespread adoption of VLMs in various applications brings significant safety challenges, particularly due to inherited vulnerabilities from traditional LLMs, such as the susceptibility to jailbreak attacks Carlini et al.
+A typical MLLM architecture is illustrated in Figure 1.
+Figure 1: The architecture of a typical MLLM.
+The widespread adoption of MLLMs in various applications brings significant safety challenges, particularly due to inherited vulnerabilities from traditional LLMs, such as the susceptibility to jailbreak attacks Carlini et al.
 (2024); Li et al.
@@
 (2024).
-Jailbreak attacks refer to an exploitative technique where malicious users can craft sophisticated-designed prompts to lead LLMs to answer misleading or harmful questions, effectively breaking the alignment and bypassing the model’s safeguard.
-Various jailbreak attack algorithms targeting LLMs have been proposed, which can be categorized into template-based Deng et al.
+Jailbreak attacks refer to an exploitative technique where malicious users can craft sophisticated-designed prompts to lead LLMs to answer misleading or harmful questions, effectively breaking the safety alignment.
+Various jailbreak attack algorithms targeting LLMs have been proposed, which can mainly be categorized into template-based Deng et al.
 (2024); Chao et al.
 (2023); Li et al.
-(2023) and optimize-based Zou et al.
+(2023) and optimization-based Zou et al.
 (2023) approaches.
-Additionally, VLMs not only inherit the vulnerabilities of LLMs but also become more susceptible to jailbreak attacks due to their integration of the visual modality.
-On the one hand, jailbreak attacks against VLMs can originate from both the textual and visual modalities, significantly broadening the scope of potential adversarial examples Gong et al.
+Additionally, MLLMs not only inherit the vulnerabilities of LLMs but also become more susceptible to jailbreak attacks due to their integration with other modalities.
+On the one hand, jailbreak attacks against MLLMs can originate from both the textual and visual modalities, significantly broadening the scope of potential adversarial examples Gong et al.
 (2023); Shayegani et al.
 (2023).
-On the other hand, recent research indicates that fine-tuning VLMs to learn the vision modality can cause LLMs to disregard their previously learned safety alignment Zong et al.
+On the other hand, recent research indicates that fine-tuning MLLMs to learn the vision modality can cause LLMs to disregard their previously learned safety alignment Zong et al.
 (2024).
-Figure 2: The workflow of safeguarding VLM against jailbreak attacks via CIDER.
-The existing jailbreak attacks on VLMs can be categorized into two strategies.
+Figure 2: The workflow of safeguarding MLLM against jailbreak attacks via CIDER.
+The existing jailbreak attacks on MLLMs can be categorized into two strategies.
 One is white-box optimization-based attacks, which define a loss function to generate imperceptible perturbations in the image modality Carlini et al.
@@
 In addition, Zong et al.
-(2024) creates a vision-language dataset named VLGuard containing both safe and unsafe queries and images, which can be used to fine-tune VLMs for improved safety against jailbreak attacks.
+(2024) creates a vision-language dataset named VLGuard containing both safe and unsafe queries and images, which can be used to fine-tune MLLMs for improved safety against jailbreak attacks.
 However, the effectiveness of VLGuard is only tested on FigStep attack and it requires the model to be white-box to fine-tune.
@@
 (2023) proposed a mutation-based jailbreaking detection framework named Jailguard.
-However, the performance of Jailguard heavily relies on the VLMs’ original safety alignment, and it significantly increases computational costs during the inference phase.
+However, the performance of Jailguard heavily relies on the MLLMs’ original safety alignment, and it significantly increases computational costs during the inference phase.
 Multimodal information is a double-edged sword: while it increases the risk of attacks, it also provides additional data that helps enhance safeguards.
 Inspired by this potential, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, specifically targeting optimization-based jailbreak attacks that are more imperceptible and susceptible.
-The intuition is that optimization-based perturbations break the VLM’s safeguards by capturing the main harmful content in the malicious query.
+The intuition is that optimization-based perturbations break the MLLM’s safeguards by capturing the main harmful content in the malicious query.
 As a result, the semantic distance between a harmful query and an adversarially perturbed image is significantly smaller than that between a harmful query and a clean image.
@@
 However, the adversarially perturbed image designed for jailbreak will experience a significant drop.
-By setting a threshold based on this change, we can effectively detect adversarially perturbed images aimed at jailbreaking VLMs.
+By setting a threshold based on this change, we can effectively detect adversarially perturbed images aimed at jailbreaking MLLMs.
 The detailed intuition is elaborated in Section 2.
-As a pre-detection module encapsulated before any VLMs, the key advantage of CIDER is its plug-and-play nature, making it independent of the target model.
-Additionally, timely inference is crucial for safeguarding VLMs.
-CIDER achieves this by adding only denoiser procedures, ensuring efficient without introducing significant inference latency.
+As a pre-detection module encapsulated before any MLLMs, the key advantage of CIDER is its plug-and-play nature, making it independent of the target model.
+Additionally, timely inference is crucial for safeguarding MLLMs.
+CIDER achieves this with denoising procedures, ensuring efficiency and minimal latency.
 In this work, we propose CIDER, an effective and efficient pre-detection module that denoises and inspects each input image.
-For images identified as adversarially perturbed for jailbreak purposes (where the semantic shift exceeds a pre-defined threshold), the VLM will refuse to generate a response.
-Images deemed normal will be processed along with the text input for model inference by the VLM.
-The workflow of safeguarding VLMs against jailbreak attacks using CIDER is illustrated in Figure 2.
+For images identified as adversarially perturbed for jailbreak purposes (where the semantic shift exceeds a pre-defined threshold), the MLLM will refuse to generate a response.
+Images deemed normal will be processed along with the text input for model inference by the MLLM.
+The workflow of safeguarding MLLMs against jailbreak attacks using CIDER is illustrated in Figure 2.
 Our contribution can be summarized as follows:
@@
 By incorporating a diffusion-based denoiser to uncover the potential of mitigating harmful information in adversarial images through denoising.
-  • We propose a plug-and-play jailbreaking detector, CIDER, which can effectively safeguard VLMs while incurring almost no additional computational overhead.
+  • We propose a plug-and-play jailbreaking detector, CIDER, which can effectively safeguard MLLMs while incurring almost no additional computational overhead.
   • Extensive experiments validate that CIDER outperforms the baseline method, achieving a higher detection success rate while reducing the computational cost as well.
-Experimental results also demonstrate its strong transferability across both white-box and black-box VLMs and attack methods.
+Experimental results also demonstrate its strong transferability across both white-box and black-box MLLMs and attack methods.
 2 Intuition: Cross-modality information is a double-edged sword
 While multimodal information aggravates model vulnerability to jailbreak attacks, it also provides additional information for defense.
-The design of CIDER is based on the intuition that optimization-based jailbreak attacks break the VLM’s safeguards by sharing harmful content in the malicious query to the image modality.
+The design of CIDER is based on the intuition that optimization-based jailbreak attacks break the MLLM’s safeguards by sharing harmful content in the malicious query to the image modality.
 Consequently, the adversarially perturbed image is closer to the harmful query in the semantic space than the clean images.
-To support this intuition, we first explain the fundamentals of the optimization-based jailbreak attacks on VLMs.
-Then, we design a few experiments to explore how cross-modal analysis can help safeguard VLMs, and we analyze the semantic difference between clean and adversarial images relative to harmful queries, both before and after denoising.
-2.1 Preliminaries: Optimization-based Jailbreak Attacks on VLMs
-Optimization-based VLM jailbreaking is similar to adversarial attacks on image classification tasks Goodfellow et al.
+To support this intuition, we first explain the fundamentals of the optimization-based jailbreak attacks on MLLMs.
+Then, we design a few experiments to explore how cross-modal analysis can help safeguard MLLMs, and we analyze the semantic difference between clean and adversarial images relative to harmful queries, both before and after denoising.
+2.1 Preliminaries: Optimization-based Jailbreak Attacks on MLLMs
+Optimization-based MLLM jailbreaking is similar to adversarial attacks on image classification tasks Goodfellow et al.
 (2014), with the primary difference being the difference in the loss function.
-Specifically, given a dataset $D=\{(q,a)\}$ where $q$ represents the harmful queries and $a$ is the corresponding targeted answers, the attacker aims to find an adversarial image $x_{adv}$ that can encourage the VLM $\mathcal{F}$ to generate $a$ when inputting $q$ along with $x_{adv}$.
+Specifically, given a dataset $D=\{(q,a)\}$ where $q$ represents the harmful queries and $a$ is the corresponding targeted answers, the attacker aims to find an adversarial image $x_{adv}$ that can encourage the MLLM $\mathcal{F}$ to generate $a$ when inputting $q$ along with $x_{adv}$.
 The objective can be formulated as:
 $x_{adv}=\underset{x_{adv}\in[0,1]^{d}}{\mathrm{argmin}}log(\mathcal{F}(a|q,x_{adv}))\vspace{-.5em}$ (1)
-where $\mathcal{F}(a|q,x_{adv})$ represents the likelihood that the VLM $\mathcal{F}$ generate answer $a$ when given the adversarial image $x_{adv}$ and the query $q$.
+where $\mathcal{F}(a|q,x_{adv})$ represents the likelihood that the MLLM $\mathcal{F}$ generate answer $a$ when given the adversarial image $x_{adv}$ and the query $q$.
 2.2 Experimental Setup
-We design a series of experiments to explore how cross-modality information can help safeguard VLMs and to analyze the semantic difference between clean and adversarial images to harmful queries, before and after denoising.
-We utilize the image and text encoder of the state-of-the-art VLM LLaVA-v1.5-7B Liu et al.
+We design a series of experiments to explore how cross-modality information can help safeguard MLLMs and to analyze the semantic difference between clean and adversarial images to harmful queries, before and after denoising.
+We utilize the image and text encoder of the state-of-the-art MLLM LLaVA-v1.5-7B Liu et al.
 (2024) to capture the semantic meanings.
@@
 In terms of denoiser, we incorporate a diffusion-based denoiser Nichol and Dhariwal (2021) to preprocess the image modality.
-The inputs to the VLMs consist of two modalities: images and text queries.
+The inputs to the MLLMs consist of two modalities: images and text queries.
 For malicious queries, we utilize the validation set proposed in the Harmbench framework Mazeika et al.
@@
 Figure 3(c) illustrates how cosine similarity between harmful query and adversarial images decreases as denoising progresses, indicating a reduction in harmful information.
-Despite this reduction, denoised adversarial images, when tested with harmful text inputs in the VLM, still enabled a significant number of queries to jailbreak.
+Despite this reduction, denoised adversarial images, when tested with harmful text inputs in the MLLM, still enabled a significant number of queries to jailbreak.
 Thus, while denoising mitigates harmful information in images, it does not eliminate their adversarial properties.
@@
 3.1 Overview
-CIDER is implemented on top of the VLMs to defense optimization-based adversarial jailbreak attacks.
+CIDER is implemented on top of the MLLMs to defense optimization-based adversarial jailbreak attacks.
 Figure 2 presents the overview of the CIDER pipeline.
@@
 […]
-Input: $img(o)$: input image, $text$: input query, $\mathcal{F}$: target VLM, $\tau$: predefined threshold.
+Input: $img(o)$: input image, $text$: input query, $\mathcal{F}$: target MLLM, $\tau$: predefined threshold.
 $flag\leftarrow true$;
@@
 We then evaluate the effectiveness and efficiency of CIDER, comparing with the baseline methods.
-Next, we discuss the trade-off between robustness and utility.
+Next, we discuss the trade-off between robustness and utility, and the choice of denoising method.
 Finally, we demonstrate the generalization of our method.
@@
 Models.
-Note that CIDER is an auxiliary model that is independent to the VLMs.
-We use LLaVA to capture the semantic meaning of each modality, but CIDER can be plugged into any other VLMs.
-To demonstrate the effectiveness of CIDER, we test the detection and defense performance on four open-source VLMs, including LLaVA-v1.5-7B Liu et al.
+Note that CIDER is an auxiliary model that is independent of the MLLMs.
+We use LLaVA to capture the semantic meaning of each modality, but CIDER can be plugged into any other MLLMs.
+To demonstrate the effectiveness of CIDER, we test the detection and defense performance on four open-source MLLMs, including LLaVA-v1.5-7B Liu et al.
 (2024), MiniGPT4 Zhu et al.
@@
 (2024), and Qwen-VL Bai et al.
-(2023), as well as the API-access VLM, GPT4V Achiam et al.
+(2023), as well as the API-access MLLM, GPT4V Achiam et al.
 (2023).
-Figure 5: ASR of base VLM, defending with CIDER and defending with Jailguard
+Figure 5: ASR of base MLLM, defending with CIDER and defending with Jailguard
 Datasets.
@@
 (2024).
-To further demonstrate CIDER will not destroy the original utilities on the normal queries, we also evaluate the utility of CIDER protected VLMs on MM-Vet benchmark Yu et al.
+To further demonstrate CIDER will not destroy the original utilities on the normal queries, we also evaluate the utility of CIDER protected MLLMs on MM-Vet benchmark Yu et al.
 (2023), which examines 6 core vision language capabilities, including recognition, optical character recognition (OCR), knowledge, language generation, spatial awareness, and math.
@@
 We use Jailguard Zhang et al.
-(2023) as a baseline, which is a SoTA mutation-based jailbreak detection strategy that protects the VLMs at the inference stage.
+(2023) as a baseline, which is a SoTA mutation-based jailbreak detection strategy that protects the MLLMs at the inference stage.
 We involve four evaluation metrics to demonstrate the performance of defending methods from different aspects.
@@
 […]
-ASR is a standard evaluation metric indicating the proportion of samples that can successfully jailbreak VLM $\mathcal{F}$ and generate harmful contents, which can be stated as:
+ASR is a standard evaluation metric indicating the proportion of samples that can successfully jailbreak MLLM $\mathcal{F}$ and generate harmful contents, which can be stated as:
 […]
@@
 (2024) that determines the harmfulness of a response.
-$\mathbb{I}_{harm}$ and $\mathbb{I}_{harm}$ represent the adversarial and harmful indicator.
+$\mathbb{I}_{adv}$ and $\mathbb{I}_{harm}$ represent the adversarial and harmful indicator.
 In terms of efficiency, we measure the time cost of processing a single query.
-In addition, to evaluate the model utility on regular tasks, responses, we incorporate an online evaluator MM-Vet-Evaluator (2024) provided along with MM-Vet benchmark, which utilizes GPT-4 to generate a soft grading score from 0 to 1 for each answer.
+In addition, to evaluate the model utility on regular tasks, and responses, we incorporate an online evaluator MM-Vet-Evaluator (2024) provided along with MM-Vet benchmark, which utilizes GPT-4 to generate a soft grading score from 0 to 1 for each answer.
 4.2 Effectiveness
@@
 We first demonstrate the overall DSR that CIDER can achieve and compare it with the baseline method, Jailguard.
-Table 1 shows that CIDER achieves a DSR of approximately 80%, while the DSR of Jailguard varies, depending on the target VLMs.
-Note that CIDER is independent of the VLMs, thus the DSR does not vary with the choice of VLMs.
+Table 1 shows that CIDER achieves a DSR of approximately 80%, while the DSR of Jailguard varies, depending on the target MLLMs.
+Note that CIDER is independent of the MLLMs, thus the DSR does not vary with the choice of MLLMs.
 However, Jailguard’s detection capability relies heavily on the model’s safety alignment, so the DSR also varies.
-VLMs with good alignment achieve high DSR (e.g., GPT4V), while poorly aligned VLMs have relatively low DSR (e.g., InstructBLIP).
-In other words, Jailguard does not significantly enhance VLM robustness against adversarial jailbreak attacks, whereas CIDER does.
+MLLMs with good alignment achieve high DSR (e.g., GPT4V), while poorly aligned MLLMs have relatively low DSR (e.g., InstructBLIP).
+In other words, Jailguard does not significantly enhance MLLM robustness against adversarial jailbreak attacks, whereas CIDER does.
 Nonetheless, CIDER achieves a higher DSR than most of the Jailguard results, except Jailguard on GPT4V.
@@
 To evaluate the effectiveness of CIDER, we measure the decline in ASR after applying CIDER.
-Figure 5 compares the original ASR without defense (red bar), ASR after CIDER (blue bar) and ASR after Jailguard (yellow bar).
+Figure 5 compares the original ASR without defense (red bar), ASR after CIDER (blue bar), and ASR after Jailguard (yellow bar).
 Note that, Jailguard is solely designed to detect jailbreak input.
@@
 Specifically, if Jailguard detects a jailbreak, it will refuse to respond, similar to CIDER.
-Otherwise, the original input will be processed by the VLM.
+Otherwise, the MLLM will process the original input.
 Across all models, defending with CIDER significantly reduces the ASR, yielding better results than the baseline.
-This indicates that CIDER effectively enhances the robustness of VLMs against optimization-based jailbreak attacks.
+This indicates that CIDER effectively enhances the robustness of MLLMs against optimization-based jailbreak attacks.
 The most notable improvements are seen in LLaVA-v1.5-7B, where ASR drops from 60% to 0%, and in MiniGPT4, from 57% to 9%.
-For VLMs with initially low ASRs, such as InstructBLIP and Qwen-VL, ASR is reduced to approximately 2% and 1% respectively.
+For MLLMs with initially low ASRs, such as InstructBLIP and Qwen-VL, ASR is reduced to approximately 2% and 1% respectively.
 Another notable disadvantage of Jailguard is observed in models like GPT4V, InstructBLIP, and Qwen-VL, which already had strong safety alignment and resistance to adversarial attacks.
 In these cases, the use of Jailguard resulted in a slight increase in ASR.
+We conclude that the threshold determined by CIDER can be effectively applied to different MLLMs due to their shared transformer-based LLM backbones, which generate comparable representations of harmful information.
+This harmful information, distilled from malicious queries, is embedded into adversarial images using similar optimization-based attacks.
+As a result, the consistent noise patterns produced by these attacks across different MLLMs can be detected using the same threshold, highlighting the robustness and transferability of CIDER.
 4.3 Efficiency
-Timely inference is crucial for safeguarding VLMs in real-world applications.
-Table 2 shows the time required to process a single input pair and generate up to 300 tokens with different VLMs, comparing no defense, CIDER, and Jailguard.
+Timely inference is crucial for safeguarding MLLMs in real-world applications.
+Table 2 shows the time required to process a single input pair and generate up to 300 tokens with different MLLMs, comparing no defense, CIDER, and Jailguard.
 Model Original CIDER Jailguard
@@
 4.4 Robustness-utility trade-off
-Figure 6: VLM performance with and without CIDER on MM-Vet.
-To further demonstrate CIDER’s influence on the original utilities on normal queries, we also evaluate the utility of CIDER protected VLMs on MM-Vet benchmark, including recognition, OCR, knowledge, language generation, spatial awareness, and math.
+Figure 6: MLLM performance with and without CIDER on MM-Vet.
+To further demonstrate CIDER’s influence on the original utilities on normal queries, we also evaluate the utility of CIDER protected MLLMs on MM-Vet benchmark, including recognition, OCR, knowledge, language generation, spatial awareness, and math.
 As shown in Figure 6, employing CIDER leads to an approximate 30% overall performance decline on normal tasks.
-Specifically, CIDER mostly affects the VLM’s recognition, knowledge, and language generation capabilities, while it has minimal impact on OCR, spatial awareness, and math skills.
+Specifically, CIDER mostly affects the MLLM’s recognition, knowledge, and language generation capabilities, while it has minimal impact on OCR, spatial awareness, and math skills.
 We hypothesize that CIDER’s stringent decision-making process, which outright rejects tasks once an image is identified as adversarial, hampers the model’s overall performance.
@@
 This approach could reduce the negative impact on the model’s functionality and we leave it to our future work.
+4.5 Ablation study on denoising method
+We perform an ablation study on the choice of denoising method in the CIDER architecture, as it significantly impacts both defense effectiveness and general task performance.
+Image smoothing methods commonly fall into two categories: DNN-based approaches, such as diffusion-based denoisers Nichol and Dhariwal (2021), and traditional filtering methods, like the Non-local Means filter (NLM; Buades et al., 2011).
+We compare the ASR and MM-Vet scores of these two denoising methods, with the results presented in Table 3.
+The NLM filter performs similarly to the diffusion-based denoiser in terms of ASR, except on the LLaVA model, where it falls significantly behind.
+However, for general tasks, the NLM filter underperforms the diffusion-based denoiser across all models.
+This suggests that while the diffusion-based denoiser effectively reduces noise, it also preserves essential image details, making it a superior choice overall.
+[…]
+Denoiser NLM Denoiser NLM
+LLaVA-v1.5-7B $0.00$ $8.12$ $19.1$ $18.9$
+InstructBLIP $2.34$ $2.56$ $12.7$ $11.1$
+MiniGPT4 $9.22$ $9.06$ $13.0$ $11.2$
+Qwen-VL $1.09$ $0.78$ $29.2$ $27.6$
+Average 3.16 $5.13$ 18.5 $17.2$
+Table 3: The ASR and MM-Vet score of different denoising methods.
 4.6 Generalization
-In the previous sections, we evaluated the ASR and DSR against adversarial examples generated by Qi et al.
+In the previous sections, we evaluated the DSR and ASR against adversarial examples generated by Qi et al.
 (2024).
-To assess the generalization of our defense method, which is crucial for its applicability to other attacks, we evaluate CIDER against another optimization-based jailbreak attack.
-We generated 800 adversarial pairs using ImgJP, as proposed by Niu et al.
+To further assess the generalization of our defense method, which is critical for its applicability to other types of attacks, we test CIDER under different attack settings.
+Dataset.
+In addition to Harmbench, we employ RedTeam-2k Luo et al.
+(2024) as a harmful query dataset, which compiles harmful queries from five different datasets and filters them according to the safety policies of OpenAI and Llama2, resulting in 2,000 red-teaming questions.
+We randomly select 200 queries and generate 800 text-image pairs using a similar processing method described in Section 2.2.
+On this dataset, CIDER achieves a DSR of 81.37%, with the reduction in ASRs presented in Table 4.
+Model Base CIDER
+ASR(%) ASR(%) $\Delta$ (%)
+LLaVA-v1.5-7B $29.87$ $1.87$ $28.00$
+InstructBLIP $24.13$ $4.37$ $19.76$
+MiniGPT4 $43.75$ $15.63$ $28.12$
+Qwen-VL $10.62$ $2.37$ $8.25$
+Table 4: Generalization to RedTeam-2k dataset
+Attack Method.
+To evaluate the generalization of our defense against different attack methods, we generated 800 adversarial pairs using ImgJP, an optimization-based approach proposed by Niu et al.
 (2024).
-Table 3 presents the drop in ASRs on four open-source VLMs.
-The ASR for all VLMs dropped to below 4%, with Qwen reaching 0%.
-Additionally, CIDER achieves a DSR of 93.87% against ImgJP.
-These results demonstrate that CIDER effectively generalizes in defending against optimization-based adversarial attacks, highlighting its practical utility for real-world applications.
-Base CIDER
-Model ASR(%) ASR(%) $\Delta$ (%)
+Table 5 shows the decrease in ASRs across four open-source MLLMs, with all ASRs falling below 4%, and Qwen-VL achieving a 0% ASR.
+Moreover, CIDER attained a DSR of 93.87% against ImgJP.
+These results demonstrate that CIDER effectively generalizes to defend against optimization-based adversarial attacks, underscoring its practical value in real-world applications.
+Model Base CIDER
+ASR(%) ASR(%) $\Delta$ (%)
 LLaVA-v1.5-7B $61.0$ $3.5$ $57.5$
@@
 Qwen-VL $6.5$ $0.0$ $6.5$
-Table 3: Generalization of CIDER to ImgJP
+Table 5: Generalization to ImgJP attack
 5 Related Work
-Vision Language Model.
-A typical Vision Language Model (VLM) consists of an image encoder Dosovitskiy et al.
+Multimodal Large Language Model.
+A typical Multimodal Large Language Model (MLLM) consists of an image encoder Dosovitskiy et al.
 (2020) to extract feature maps, a projector to align image modality information with text modality, and a Large Language Model (LLM) to integrate textual and visual input for generating responses.
@@
 (2023).
-Jailbreaking VLMs.
-Incorporating visual information into the LLM framework significantly broadens its range of applications but also introduces new security vulnerabilities, complicating the security issues of VLMs.
-Besides transferring text jailbreak templates from LLMs to VLMs Luo et al.
-(2024), effective strategies for jailbreaking VLMs include using gradient-based methods to generate adversarial images Carlini et al.
+Jailbreaking MLLMs.
+Incorporating visual information into the LLM framework significantly broadens its range of applications but also introduces new security vulnerabilities, complicating the security issues of MLLMs.
+Besides transferring text jailbreak templates from LLMs to MLLMs Luo et al.
+(2024), effective strategies for jailbreaking MLLMs include using gradient-based methods to generate adversarial images Carlini et al.
+(2024); Zhao et al.
 (2024); Qi et al.
@@
 (2023).
-This paper focuses on safeguarding VLMs against gradient-based adversarial image attacks, aiming to fortify VLMs against such sophisticated threats and ensure their robustness and reliability in practical applications.
-Safeguarding VLMs.
-Various defense mechanisms have been proposed to address vulnerabilities in VLMs and enhance their security and robustness.
+This paper focuses on safeguarding MLLMs against gradient-based jailbreak attacks via adversarial images, aiming to fortify MLLMs against such sophisticated threats and ensure their robustness and reliability in practical applications.
+Safeguarding MLLMs.
+Various defense mechanisms have been proposed to address vulnerabilities in MLLMs and enhance their security and robustness.
 These mechanisms can be categorized into proactive and reactive defenses based on their preventive and responsive nature.
@@
 Compared to previous methods, CIDER achieves superior defense performance, as evidenced by higher DSR and a significant decline in ASR, while greatly reducing processing time.
-We also evaluate the transferability of CIDER to other optimization-based adversarial attacks and demonstrate the robustness-utility trade-off in VLMs.
-In future research, we aim to improve CIDER by reducing the negative impact on VLM utilities to normal tasks.
+We also evaluate the generalization of CIDER to other datasets and optimization-based adversarial attacks, and demonstrate the robustness-utility trade-off in MLLMs.
+In future research, we aim to improve CIDER by reducing the negative impact on MLLM utilities to normal tasks.
 Additionally, it would be useful to develop defense mechanisms against non-optimization-based jailbreak attacks.
+Acknowledgement
+We thank all reviewers for their constructive comments.
+This work is supported by the Shanghai Engineering Research Center of Intelligent Vision and Imaging and the Open Research Fund of The State Key Laboratory of Blockchain and Data Security, Zhejiang University.
 Limitations
@@
 1.
-While CIDER is an effective, efficient, and user-friendly defense mechanism, it does impact VLM performance to some extent.
+While CIDER is an effective, efficient, and user-friendly defense mechanism, it does impact MLLM performance to some extent.
 We believe this is due to CIDER’s stringent handling of adversarial examples.
-In future work, we plan to implement multi-level thresholds to process adversarial examples with varying degrees of rigor, aiming to maintain robust defense capabilities without compromising VLM performance.
+In future work, we plan to implement multi-level thresholds to process adversarial examples with varying degrees of rigor, aiming to maintain robust defense capabilities without compromising MLLM performance.
 2.
 CIDER is specifically designed to defend against optimization-based adversarial jailbreak attacks, and its effectiveness against other types of jailbreak attacks is uncertain.
-Future research will explore CIDER’s effectiveness against these alternative attacks and develop corresponding defense strategies, aiming to enhance the overall security and resilience of VLMs against a wider array of adversarial threats.
+Future research will explore CIDER’s effectiveness against these alternative attacks and develop corresponding defense strategies, aiming to enhance the overall security and resilience of MLLMs against a wider array of adversarial threats.
 Ethics Statement
-Ensuring the security of Vision Large Language Models (VLMs) is crucial as they become more widely used in various applications.
-This paper introduces CIDER, a simple yet effective cross-modality information detector designed to defend against adversarial jailbreak attacks in VLMs.
+Ensuring the security of Vision Large Language Models (MLLMs) is crucial as they become more widely used in various applications.
+This paper introduces CIDER, a simple yet effective cross-modality information detector designed to defend against adversarial jailbreak attacks in MLLMs.
 Our work significantly contributes to the field by providing a tool that mitigates known vulnerabilities and lays the groundwork for future improvements in safety measures.
-While CIDER marks significant progress, it doesn’t make VLMs immune to all threats.
-Continuous evaluation and updates are crucial as VLMs evolve.
-By sharing CIDER and our findings, we aim to encourage ongoing research and collaboration, promoting advanced and secure VLMs.
-We are committed to addressing the ethical implications of VLM deployment, ensuring transparency, and prioritizing the responsible use of these technologies for societal benefit.
+While CIDER marks significant progress, it doesn’t make MLLMs immune to all threats.
+Continuous evaluation and updates are crucial as MLLMs evolve.
+By sharing CIDER and our findings, we aim to encourage ongoing research and collaboration, promoting advanced and secure MLLMs.
+We are committed to addressing the ethical implications of MLLM deployment, ensuring transparency, and prioritizing the responsible use of these technologies for societal benefit.
 Appendix A Visualization of denoising
@@
 Figure 8 shows the ASR of CIDER-de and Figure 9 shows the MM-Vet score of it.
-It can be observed that using CIDER-de hardly impacts the utility of the VLM.
+It can be observed that using CIDER-de hardly impacts the utility of the MLLM.
 However, this comes at the expense of greatly diminished defensive effectiveness.
-Figure 8: ASR of base VLM, defending with CIDER-de, CIDER and Jailguard
-Figure 9: MM-Vet score of base VLM, defending with CIDER-de, CIDER and Jailguard
+Figure 8: ASR of base MLLM, defending with CIDER-de, CIDER and Jailguard
+Figure 9: MM-Vet score of base MLLM, defending with CIDER-de and CIDER